Find the period
cos(7x)

Answer by lwsshak3(11628)

Find the period
cos(7x)


Standard form for the cosine function: y=Acos(Bx-C), with A=amplitude, 2π/B=period, and C/B=Phase shift
For given function, cos(7x), B=7
so period=2π/7
Amplitude=1, and phase shift=0
ans:
Period of given function=2π/7